# Rejection-Sampled Linear Codes for Lossy Compression and Channel Simulation

### v1

@@ VI Greedy Rejection-Sampled Syndrome Encoder
In this section, we will give an algorithm for the construction of an RSSE for the exact simulation of an additive exchangeable noise channel with noise distribution $p_{\mathbf{Z}}(\mathbf{z})$, called greedy rejection-sampled syndrome encoder (GRSSE). It is inspired by the greedy rejection sampling scheme [4], which maximizes the probability of acceptance at iteration $1$, then maximizes the probability of acceptance at iteration $2$, and so on. Unlike greedy rejection sampling, we do not have a fixed proposal distribution, but rather has the flexibility to design $\kappa_{i}$ which helps us further maximize the probability of acceptance.
@@
We now describe the computation of the argmin in (2). There are various symmetries in (2) that allow us to reduce the dimension of the problem. First, due to the random permutation $\boldsymbol{\Pi}_{i}$, the distributions $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ are always exchangeable, and hence we only need to keep track of the distribution of the type $\mathbf{p}=\mathrm{type}(\mathbf{z})$, i.e., we consider $p_{\mathbf{P}}^{(i)}(\mathbf{p})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{p}}p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ for the distribution of $\mathbf{P}=\mathrm{type}(\mathbf{Z})\in\mathcal{P}_{n}(\mathbb{F}_{q})$ when $\mathbf{Z}\sim p_{\mathbf{Z}}^{(i)}$, and consider a conditional distribution $\tilde{\kappa}_{i}(\mathbf{p}|\mathbf{s})$ from the syndrome to the type instead of $\kappa_{i}(\mathbf{z}|\mathbf{s})$. This significantly reduces the dimension of the problem since $|\mathcal{P}_{n}(\mathbb{F}_{q})|=O(n^{q-1})\ll|\mathbb{F}_{q}^{n}|=q^{n}$. Second, the constraint in Definition 3 requires $\tilde{\kappa}_{i}(\mathbf{p}|\mathbf{s})>0$ only if ther exists $\mathbf{z}$ with $\mathrm{type}(\mathbf{z})=\mathbf{p}$ and $\mathbf{z}\mathbf{H}_{i}^{\top}=\mathbf{s}$. Define the type set of the syndrome $\mathbf{s}$ to be
$t_{\mathbf{H}}(\mathbf{s})=\big{\{}\mathrm{type}(\mathbf{z}):\,\mathbf{z}\in\mathbb{F}_{q}^{n},\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}\big{\}}\in 2^{\mathcal{P}_{n}(\mathbb{F}_{q})}.$
@@
$p_{t_{\mathbf{H}}}(t):=q^{-(n-k)}|\{\mathbf{s}:\,t_{\mathbf{H}}(\mathbf{s})=t\}|$ (3)
over $t\in t_{\mathbf{H}}(\mathbb{F}_{q}^{n-k})$, i.e., it is the distribution of $T=t_{\mathbf{H}}(\mathbf{S})$ where $\mathbf{S}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k})$. For $t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$, $\mathbf{p}\in\mathcal{P}_{n}(\mathbb{F}_{q})$, let
$\gamma_{i,t,\mathbf{p}}:=q^{-(n-k)}\sum_{\mathbf{s},\mathbf{z}:t_{\mathbf{H}}(\mathbf{s})=t,\,\mathrm{type}(\mathbf{z})=\mathbf{p}}\kappa_{i}(\mathbf{z}|\mathbf{s}).$
Note that $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$ is generally not $1$ since the random mapping $\kappa_{i}$ may have a positive probability of mapping $\mathbf{s}$ to the erasure symbol $\mathrm{e}$ which indicates a rejection. We can maximize the probability of acceptance by maximizing $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$. Hence, the problem (2) is equivalent to the problem of maximizing $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$ subject to the linear constraints
$\begin{array}[]{ll}\gamma_{i,t,\mathbf{p}}\geq 0,\;\forall t,\mathbf{p};&\sum_{\mathbf{p}}\gamma_{i,t,\mathbf{p}}\leq p_{t_{\mathbf{H}_{i}}}(t),\;\forall t;\\ \sum_{t}\gamma_{i,t,\mathbf{p}}\leq p_{\mathbf{P}}^{(i)}(\mathbf{p}),\;\forall\mathbf{p};&\gamma_{i,t,\mathbf{p}}=0\;\;\text{if}\;\mathbf{p}\notin t.\end{array}$
@@
[…]
    Input: type distribution $p_{\mathbf{P}}(\mathbf{p})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{p}}p_{\mathbf{Z}}(\mathbf{z})$,
@@
1:$p_{\mathbf{P}}^{(1)}(\mathbf{p})\leftarrow p_{\mathbf{P}}(\mathbf{p})$
2:$S\leftarrow 1$
@@
4:     $p_{t_{\mathbf{H}_{i}}}(t)\leftarrow q^{-(n-k)}|\{\mathbf{s}:\,t_{\mathbf{H}}(\mathbf{s})=t\}|$, $\forall t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$
5:     Construct weighted directed graph $G$ with
6:        vertices $V=\{\mathrm{src},\mathrm{des}\}\cup t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})\cup\mathcal{P}_{n}(\mathbb{F}_{q})$,
7:        edges $(t,\mathbf{p})$ with capacity $\infty$ for $t,\mathbf{p}$ with $\mathbf{p}\in t$,
8:        edges $(\mathrm{src},t)$ with cap. $p_{t_{\mathbf{H}_{i}}}(t)$ for $t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$,
9:        edges $(\mathbf{p},\mathrm{des})$ with cap. $p_{\mathbf{P}}^{(i)}(\mathbf{p})$ for $\mathbf{p}\in\mathcal{P}_{n}(\mathbb{F}_{q})$
10:     Compute maximum flow of $G$ from $\mathrm{src}$ to $\mathrm{des}$
11:     $F\leftarrow$ total flow, $\;\gamma_{i,t,\mathbf{p}}\leftarrow$ flow along edge $(t,\mathbf{p})$
12:     $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow 0$, $\forall\mathbf{s},\mathbf{z}$
13:     $\mathrm{vis}(\mathbf{s},\mathbf{p})\leftarrow 0$, $\forall\mathbf{s},\mathbf{p}$
14:     for $\mathbf{z}\in\mathbb{F}_{q}^{n}$ do
15:         $\mathbf{s}\leftarrow\mathbf{z}\mathbf{H}_{i}^{\top}$, $\;t\leftarrow t_{\mathbf{H}_{i}}(\mathbf{s}),$ $\;\mathbf{p}\leftarrow\mathrm{type}(\mathbf{z})$
16:         if $\mathrm{vis}(\mathbf{s},\mathbf{p})=0$ then
17:              $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow\gamma_{i,t,\mathbf{p}}/p_{t_{\mathbf{H}_{i}}}(t)$, $\;\mathrm{vis}(\mathbf{s},\mathbf{p})\leftarrow 1$
18:         end if
19:     end for
20:     $\kappa_{i}(\mathrm{e}|\mathbf{s})\leftarrow 1-\sum_{\mathbf{p}\in t}\gamma_{i,t,\mathbf{p}}/p_{t_{\mathbf{H}_{i}}}(t)$ (where $t=t_{\mathbf{H}_{i}}(\mathbf{s})$)
21:     $p_{L}(i)\leftarrow SF$
22:     if $F=1$ then
23:         return $(\mathbf{H}_{j},\kappa_{j})_{j\in[i]}$, $p_{L}$
24:     end if
25:     $p_{\mathbf{P}}^{(i+1)}(\mathbf{p})\leftarrow(p_{\mathbf{P}}^{(i)}(\mathbf{p})-\sum_{\mathbf{s}}\gamma_{i,\mathbf{s},\mathbf{p}})/(1-F)$
26:     $S\leftarrow S(1-F)$
27:end for
@@
VII Theoretical Analysis of GRSSE
We now analyze the performance of GRSSE. For simplicity, we consider the inner codes $\mathbf{H}_{i}=\mathbf{H}$ to be the same for all $i$. We call this the $\mathbf{H}$-GRSSE or pure GRSSE. Let the distance of the code be $d:=\min_{\mathbf{z}\neq 0:\,\mathbf{z}\mathbf{H}^{\top}=0}\mathrm{wt}(\mathbf{z})$. Theorem 6 shows that $\mathbf{H}$-GRSSE achieves the capacity under very mild conditions. Note that the bound in Theorem 6 holds regardless of the distribution of the input $\mathbf{X}$. A proof sketch will be presented later in this section. The complete proof is in Appendix -B.
@@ VII Theoretical Analysis of GRSSE
The $\mathbf{H}$-GRSSE for the simulation of the additive exchangeable noise channel with noise distribution $p_{\mathbf{Z}}$, with $\mathbf{H}_{i}=\mathbf{H}$ where the code $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ has a distance $d$, gives a number of iteration $L$ satisfying
@@
To achieve the capacity $C$, we need $d$ to be large enough so $\mathrm{wt}(\mathbf{Z})<d/2$ with high probability. A direct corollary is that for every sequence of AXN channels with noise distributions $p_{\mathbf{Z}^{(n)}}$ over $\mathbb{F}_{q}^{n}$ and capacities $C^{(n)}$ for $n=1,2,\ldots$, with asymptotic capacity $\overline{C}=\lim_{n\to\infty}C^{(n)}/n$, as long as we find a sequence of inner codes $\mathbf{H}^{(n)}$ with distance $d^{(n)}$ such that $\mathbb{P}(\mathrm{wt}(\mathbf{Z}^{(n)})\geq d^{(n)}/2)\to 0$, then GRSSE achieves the asymptotic capacity, in the sense that the expected number of bits of communication per dimension $n$ approaches $\overline{C}$. As a result, by simulating the sequence of Hamming ball channels $p_{\mathbf{Z}^{(n)}}=\mathrm{Unif}(\{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathrm{wt}(\mathbf{z})\leq\alpha n\})$ with $\overline{C}=\log q+(1-\alpha)\log(1-\alpha)+\alpha\log(\alpha/(q-1))$ for $\alpha\leq 1-1/q$, as long as the distances satisfy $\mathrm{liminf}_{n}d^{(n)}/n>2\alpha$, GRSSE can achieve the rate-distortion function for Hamming distortion.
@@
Before we present the proof of Theorem 6, we introduce a lemma which may be of interest beyond GRSSE. Unlike greedy rejection sampling (GRS) [4] which has uniquely defined and tractable iterations, the only information we know about the iterations in GRSSE is that $\kappa_{i}$ gives a low rejection probability $\mathbb{E}[\kappa_{i}(\mathrm{e}|\mathbf{S}_{i})]$ (or high acceptance probability) in (2) (the argmin might not even be unique). This flexibility poses a significant challenge to the analysis. The following lemma, which generalizes the bound for GRS in [4], shows that the expected logarithm of the number of iterations can be bounded simply by virtue of having a high acceptance probability for each iteration. It also improves upon [4, 9] by giving a smaller constant. The proof is in Appendix -A.
@@
[…]
@@
We give a sketch of the proof of Theorem 6. The complete proof is in Appendix -B. We apply Lemma 7 on $(\mathbf{Z},L)$, $\theta=1$,
$Q(\mathbf{z})=\frac{1}{2}\mathbf{1}\Big{\{}\mathrm{wt}(\mathbf{Z})<\frac{d}{2}\Big{\}}q^{k-n}+\frac{1}{2}q^{-n},$
which will yield the desired bounds. [^3] Technically $\sum_{\mathbf{z}}Q(\mathbf{z})\neq 1$. Refer to Appendix -B for the rigorous proof. To check that the acceptance probability at iteration $i$ is at least $F:=\sum_{\mathbf{z}}\min\{p_{\mathbf{Z}}^{(i)}(\mathbf{z}),\,Q(\mathbf{z})\}$ as required by Lemma 7, we need to find one particular $\kappa_{i}(\mathbf{z}|\mathbf{s})$ which gives an acceptance probability at least $F$. Given $\mathbf{s}$, this $\kappa_{i}$ either picks a $\mathbf{z}$ uniformly at random among $\mathbf{z}$’s satisfying $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$ with probability $1/2$, or picks the coset leader (the $\mathbf{z}$ giving the smallest $\mathrm{wt}(\mathbf{z})$ with $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$) with probability $1/2$. This $\kappa_{i}$ will ensure that a $\mathbf{z}$ with small $\mathrm{wt}(\mathbf{z})$ is selected with enough probability, and hence simulating the part of the noise distribution where the noise weight is small (usually the hardest part to simulate). Since a $\mathbf{z}$ with $\mathrm{wt}(\mathbf{Z})<d/2$ must be a coset leader, we can verify that the flow is indeed at least $F$.
VIII Mixed GRSSE and Experiments

### v2

@@ VII Greedy Rejection-Sampled Syndrome Encoder
In this section, we will give an algorithm for the construction of an RSSE for the exact simulation of an AXN channel with noise distribution $p_{\mathbf{Z}}(\mathbf{z})$, called greedy rejection-sampled syndrome encoder (GRSSE). It is inspired by the greedy rejection sampling scheme [4], which maximizes the probability of acceptance at iteration $1$, then maximizes the probability of acceptance at iteration $2$, and so on. Unlike greedy rejection sampling, we do not have a fixed proposal distribution, but rather has the flexibility to design $\kappa_{i}$ which helps us further maximize the probability of acceptance.
@@
We now describe the computation of the argmin in (9). There are various symmetries in (9) that allow us to reduce the dimension of the problem. First, due to the random permutation $\boldsymbol{\Pi}_{i}$, the distributions $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ are always exchangeable, and hence we only need to keep track of the distribution of the type $\mathbf{t}=\mathrm{type}(\mathbf{z})$, i.e., we consider $p_{\mathbf{T}}^{(i)}(\mathbf{t})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{t}}p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ for the distribution of $\mathbf{T}=\mathrm{type}(\mathbf{Z})\in\mathcal{P}_{n}(\mathbb{F}_{q})$ when $\mathbf{Z}\sim p_{\mathbf{Z}}^{(i)}$, and consider a conditional distribution $\tilde{\kappa}_{i}(\mathbf{t}|\mathbf{s})$ from the syndrome to the type instead of $\kappa_{i}(\mathbf{z}|\mathbf{s})$. This significantly reduces the dimension of the problem since $|\mathcal{P}_{n}(\mathbb{F}_{q})|=O(n^{q-1})\ll|\mathbb{F}_{q}^{n}|=q^{n}$. Second, the constraint in Definition 3 requires $\tilde{\kappa}_{i}(\mathbf{t}|\mathbf{s})>0$ only if there exists $\mathbf{z}$ with $\mathrm{type}(\mathbf{z})=\mathbf{t}$ and $\mathbf{z}\mathbf{H}_{i}^{\top}=\mathbf{s}$. Define the type set of the syndrome $\mathbf{s}$ to be
$\mathcal{T}_{\mathbf{H}}(\mathbf{s})=\big\{\mathrm{type}(\mathbf{z}):\,\mathbf{z}\in\mathbb{F}_{q}^{n},\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}\big\}\subseteq\mathcal{P}_{n}(\mathbb{F}_{q}).$
@@
$p_{\mathcal{T}_{\mathbf{H}}}(\mathcal{T}):=q^{-(n-k)}|\{\mathbf{s}:\,\mathcal{T}_{\mathbf{H}}(\mathbf{s})=\mathcal{T}\}|$ (10)
over $\mathcal{T}\in\mathcal{T}_{\mathbf{H}}(\mathbb{F}_{q}^{n-k})$, i.e., it is the distribution of $\mathcal{T}_{\mathbf{H}}(\mathbf{S})$ where $\mathbf{S}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k})$. For $\mathcal{T}\in\mathcal{T}_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k_{i}})$, $\mathbf{t}\in\mathcal{P}_{n}(\mathbb{F}_{q})$, let
$\gamma_{i,\mathcal{T},\mathbf{t}}:=q^{-(n-k)}\sum_{\mathbf{s},\mathbf{z}:\mathcal{T}_{\mathbf{H}}(\mathbf{s})=\mathcal{T},\,\mathrm{type}(\mathbf{z})=\mathbf{t}}\kappa_{i}(\mathbf{z}|\mathbf{s}).$
Note that $\sum_{\mathcal{T},\mathbf{t}}\gamma_{i,\mathcal{T},\mathbf{t}}$ is generally not $1$ since the random mapping $\kappa_{i}$ may have a positive probability of mapping $\mathbf{s}$ to the rejection symbol $\mathrm{e}$. We can maximize the probability of acceptance by maximizing $\sum_{\mathcal{T},\mathbf{t}}\gamma_{i,\mathcal{T},\mathbf{t}}$. Hence, the problem (9) is equivalent to the problem of maximizing $\sum_{\mathcal{T},\mathbf{t}}\gamma_{i,\mathcal{T},\mathbf{t}}$ subject to the linear constraints
$\begin{array}[]{ll}\gamma_{i,\mathcal{T},\mathbf{t}}\geq 0,\;\forall\mathcal{T},\mathbf{t};&\sum_{\mathbf{t}}\gamma_{i,\mathcal{T},\mathbf{t}}\leq p_{\mathcal{T}_{\mathbf{H}_{i}}}(\mathcal{T}),\;\forall\mathcal{T};\\ \sum_{\mathcal{T}}\gamma_{i,\mathcal{T},\mathbf{t}}\leq p_{\mathbf{T}}^{(i)}(\mathbf{t}),\;\forall\mathbf{t};&\gamma_{i,\mathcal{T},\mathbf{t}}=0\;\;\text{if}\;\mathbf{t}\notin\mathcal{T}.\end{array}$
@@
Procedure $\textsc{GRSSE\_Precompute}(p_{\mathbf{T}}(\mathbf{t}),(\mathbf{H}_{i})_{i}):$
    Input: type distribution $p_{\mathbf{T}}(\mathbf{t})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{t}}p_{\mathbf{Z}}(\mathbf{z})$,
@@
[…]
@@
4:  $p_{\mathcal{T}_{\mathbf{H}_{i}}}(\mathcal{T})\leftarrow q^{-(n-k_{i})}|\{\mathbf{s}:\,\mathcal{T}_{\mathbf{H}_{i}}(\mathbf{s})=\mathcal{T}\}|$,
5:   $\forall\mathcal{T}\in\mathcal{T}_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k_{i}})$
6:  Construct weighted directed graph $G$ with
7:   vertices $V=\{\mathrm{src},\mathrm{des}\}\cup\mathcal{T}_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k_{i}})\cup\mathcal{P}_{n}(\mathbb{F}_{q})$,
8:   edges $(\mathcal{T},\mathbf{t})$ with capacity $\infty$ for $\mathcal{T},\mathbf{t}$ with $\mathbf{t}\in\mathcal{T}$,
9:   edges $(\mathrm{src},\mathcal{T})$ with capacity $p_{\mathcal{T}_{\mathbf{H}_{i}}}(\mathcal{T})$
10:    for $\mathcal{T}\in\mathcal{T}_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k_{i}})$,
11:   edges $(\mathbf{t},\mathrm{des})$ with capacity $p_{\mathbf{T}}^{(i)}(\mathbf{t})$ for $\mathbf{t}\in\mathcal{P}_{n}(\mathbb{F}_{q})$
12:  Compute maximum flow of $G$ from $\mathrm{src}$ to $\mathrm{des}$
13:  $F\leftarrow$ total flow, $\gamma_{i,\mathcal{T},\mathbf{t}}\leftarrow$ flow along edge $(\mathcal{T},\mathbf{t})$
14:  $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow 0$, $\forall\mathbf{s},\mathbf{z}$
15:  $\mathrm{vis}(\mathbf{s},\mathbf{t})\leftarrow 0$, $\forall\mathbf{s},\mathbf{t}$
16:  for $\mathbf{z}\in\mathbb{F}_{q}^{n}$ do
17:   $\mathbf{s}\leftarrow\mathbf{z}\mathbf{H}_{i}^{\top}$, $\mathcal{T}\leftarrow\mathcal{T}_{\mathbf{H}_{i}}(\mathbf{s}),$ $\mathbf{t}\leftarrow\mathrm{type}(\mathbf{z})$
18:   if $\mathrm{vis}(\mathbf{s},\mathbf{t})=0$ then
19:     $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow\gamma_{i,\mathcal{T},\mathbf{t}}/p_{\mathcal{T}_{\mathbf{H}_{i}}}(\mathcal{T})$, $\mathrm{vis}(\mathbf{s},\mathbf{t})\leftarrow 1$
20:   end if
21:  end for
22:  $\kappa_{i}(\mathrm{e}|\mathbf{s})\leftarrow 1-\sum_{\mathbf{t}\in\mathcal{T}}\gamma_{i,\mathcal{T},\mathbf{t}}/p_{\mathcal{T}_{\mathbf{H}_{i}}}(\mathcal{T})$
23:   (where $\mathcal{T}=\mathcal{T}_{\mathbf{H}_{i}}(\mathbf{s})$)
24:  $p_{L}(i)\leftarrow SF$
25:  if $F=1$ then
26:   return $(\mathbf{H}_{j},\kappa_{j})_{j\in[i]}$, $p_{L}$
27:  end if
28:  $p_{\mathbf{T}}^{(i+1)}(\mathbf{t})\leftarrow(p_{\mathbf{T}}^{(i)}(\mathbf{t})-\sum_{\mathcal{T}}\gamma_{i,\mathcal{T},\mathbf{t}})/(1-F)$
29:  $S\leftarrow S(1-F)$
30:end for
@@
VIII Theoretical Analysis of GRSSE
We now analyze the performance of GRSSE. For simplicity, we consider the inner codes $\mathbf{H}_{i}=\mathbf{H}$ to be the same for all $i$. We call this the $\mathbf{H}$-GRSSE or pure GRSSE. Let the distance of the code be $d:=\min_{\mathbf{z}\neq 0:\,\mathbf{z}\mathbf{H}^{\top}=0}\mathrm{wt}(\mathbf{z})$. Theorem 15 shows that $\mathbf{H}$-GRSSE achieves the capacity under very mild conditions. Note that the bound in Theorem 15 holds regardless of the distribution of the input $\mathbf{X}$. A proof sketch will be presented later in this section. The complete proof is in Appendix -C.
@@ VIII Theoretical Analysis of GRSSE
The $\mathbf{H}$-GRSSE for the simulation of the AXN channel with noise distribution $p_{\mathbf{Z}}$, with $\mathbf{H}_{i}=\mathbf{H}$ where the code $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ has a distance $d$, gives a number of iteration $L$ satisfying
@@
To achieve the capacity $C$, we need $d$ to be large enough so $\mathrm{wt}(\mathbf{Z})<d/2$ with high probability. A direct corollary is that for every sequence of AXN channels with noise distributions $p_{\mathbf{Z}^{(n)}}$ over $\mathbb{F}_{q}^{n}$ and capacities $C^{(n)}$ for $n=1,2,\ldots$, with asymptotic capacity $\tilde{C}=\lim_{n\to\infty}C^{(n)}/n$, as long as we find a sequence of inner codes $\mathbf{H}^{(n)}$ with distance $d^{(n)}$ such that $\mathbb{P}(\mathrm{wt}(\mathbf{Z}^{(n)})\geq d^{(n)}/2)\to 0$, then GRSSE achieves the asymptotic capacity, in the sense that the expected number of bits of communication per dimension $n$ approaches $\tilde{C}$. As a result, by simulating the sequence of Hamming ball channels $p_{\mathbf{Z}^{(n)}}=\mathrm{Unif}(\{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathrm{wt}(\mathbf{z})\leq\alpha n\})$ with $\tilde{C}=\log q+(1-\alpha)\log(1-\alpha)+\alpha\log(\alpha/(q-1))$ for $\alpha\leq 1-1/q$, as long as the distances satisfy $\mathrm{liminf}_{n}d^{(n)}/n>2\alpha$, GRSSE can achieve the rate-distortion function for Hamming distortion.
@@
Before we present the proof of Theorem 15, we introduce a lemma which may be of interest beyond GRSSE. Unlike greedy rejection sampling (GRS) [4] which has uniquely defined and tractable iterations, the only information we know about the iterations in GRSSE is that $\kappa_{i}$ gives a low rejection probability $\mathbb{E}[\kappa_{i}(\mathrm{e}|\mathbf{S}_{i})]$ (or high acceptance probability) in (9) (the argmin might not even be unique). This flexibility poses a significant challenge to the analysis. The following lemma, which generalizes the bound for GRS in [4], shows that the expected logarithm of the number of iterations can be bounded simply by virtue of having a high acceptance probability for each iteration. It also improves upon [4, 9] by giving a smaller constant. The proof is in Appendix -B.
@@
$\mathbb{P}(L=i\,|\,L\geq i)$ $\geq\theta\int_{\mathcal{Y}}\min\left\{\frac{\mathrm{d}P_{Y|L\geq i}}{\mathrm{d}Q}(y),\,1\right\}Q(\mathrm{d}y)$ (13)
@@
We give a sketch of the proof of Theorem 15. The complete proof is in Appendix -C. We apply Lemma 16 on $(\mathbf{Z},L)$, $\theta=1$,
$Q(\mathbf{z})=\frac{1}{2}\mathbbm{1}\Big\{\mathrm{wt}(\mathbf{Z})<\frac{d}{2}\Big\}q^{k-n}+\frac{1}{2}q^{-n},$
which will yield the desired bounds. [^4] Technically $\sum_{\mathbf{z}}Q(\mathbf{z})\neq 1$. Refer to Appendix -C for the rigorous proof. To check that the acceptance probability at iteration $i$ is at least $F:=\sum_{\mathbf{z}}\min\{p_{\mathbf{Z}}^{(i)}(\mathbf{z}),\,Q(\mathbf{z})\}$ as required by Lemma 16, we need to find one particular $\kappa_{i}(\mathbf{z}|\mathbf{s})$ which gives an acceptance probability at least $F$. Given $\mathbf{s}$, this $\kappa_{i}$ either picks a $\mathbf{z}$ uniformly at random among $\mathbf{z}$’s satisfying $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$ with probability $1/2$, or picks the coset leader (the $\mathbf{z}$ giving the smallest $\mathrm{wt}(\mathbf{z})$ with $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$) with probability $1/2$. This $\kappa_{i}$ will ensure that a $\mathbf{z}$ with small $\mathrm{wt}(\mathbf{z})$ is selected with enough probability, and hence simulating the part of the noise distribution where the noise weight is small (usually the hardest part to simulate). Since a $\mathbf{z}$ with $\mathrm{wt}(\mathbf{Z})<d/2$ must be a coset leader, we can verify that the flow is indeed at least $F$.
IX Mixed GRSSE and Experiments